OpenStudy (anonymous):

Solve x2 + 8x − 3 = 0 using the completing-the-square method.

OpenStudy (anonymous):

these are the answer choices x = four plus or minus the square root of three x = negative four plus or minus the square root of three x = four plus or minus the square root of nineteen x = negative four plus or minus the square root of nineteen

OpenStudy (wolf1728):

x^2 + 8x = 3 Take the x coefficient ; divide it by 2; square it then add to both sides of the equation 8/2 = 4 4^2 = 16 x^2 + 8x +16 = 3 + 16 (X+4)^2 = 19 X+4 = sq root of 19

OpenStudy (anonymous):

wow thank you so much wolf

OpenStudy (anonymous):

Can you help me with two other questions?

OpenStudy (wolf1728):

I'm still seeing if I got that right (and also trying to see which multiple choice answer is correct).

OpenStudy (anonymous):

Using the completing-the-square method, rewrite f(x) = x2 − 8x + 3 in vertex form. f(x) = (x − 8)2 f(x) = (x − 4)2 − 13 f(x) = (x − 4)2 + 3 f(x) = (x − 4)2 + 16

OpenStudy (anonymous):

Rewrite f(x) = –(x + 3)2 − 1 from vertex form to standard form. f(x) = –x2 − 6x − 10 f(x) = x2 + 6x + 8 f(x) = x2 − 6x − 8 f(x) = x2 − 8

OpenStudy (anonymous):

please help me wolf

OpenStudy (wolf1728):

not sure about vertex form - and I'm on a phone call at the moment

OpenStudy (anonymous):

oh ok thanks for your help though

OpenStudy (whpalmer4):

Using the completing-the-square method, rewrite \(f(x) = x^2 − 8x + 3 \) in vertex form. So vertex form is just \[y = (x-h)^2 + k\]with the vertex at \((h,k)\) Complete the square and replace \(y\) with \(f(x)\) Rewrite \(f(x) =\ –(x + 3)^2 − 1\) from vertex form to standard form. Even easier — just expand the right hand side of the equation.

OpenStudy (anonymous):

thank you so much